OpenStudy (gigi1103):

Each car in a rental fleet is either red or white, and there are 75 cars total. There are three more white than twice the number of red cars. How many cars are white?

OpenStudy (gigi1103):

@beckyg111

OpenStudy (gigi1103):

@Kitten_is_back

OpenStudy (gigi1103):

@Nnesha

OpenStudy (gigi1103):

help plz

Vocaloid (vocaloid):

well, let's let r = number of red cars and w = number of white cars we can write two equations using the information in the problem r + w = 75 w = 2r + 3 know where to go from here?

OpenStudy (gigi1103):

?

Vocaloid (vocaloid):

well, ok, I'll walk you through step by step like the problem says, there are 75 total cars, so r + w = 75 does that make sense?

OpenStudy (gigi1103):

yes

Vocaloid (vocaloid):

the problem also says "there are three more white than twice the number of red cars" so w = 2r + 3 with me so far?

OpenStudy (gigi1103):

yup

Vocaloid (vocaloid):

so we put them together: r + w = 75 w = 2r + 3 now, the second equation says w = 2r + 3, so we can replace "w" with "2r+3" whenever we see w now, going back to the first equation we have: r + w = 75 r + (2r + 3) = 75 does that make sense?

OpenStudy (gigi1103):

yeah still with ya

Vocaloid (vocaloid):

ok, so we have r + (2r + 3) = 75 can you solve for r now?

OpenStudy (gigi1103):

so i divide by two

Vocaloid (vocaloid):

uh, not quite we have r + (2r + 3) = 75 let's combine the r's together r + 2r = ?

OpenStudy (gigi1103):

what do you mean?

Vocaloid (vocaloid):

just add them together r + 2r = ?

OpenStudy (gigi1103):

dang how?

OpenStudy (gigi1103):

they are numbers tho

Vocaloid (vocaloid):

r is just a letter that takes the place of a number think of it this way: one dog + two dogs = three dogs one r + 2r = 3r does that make sense?

OpenStudy (gigi1103):

yeah

OpenStudy (gigi1103):

so its 3rs now?

OpenStudy (gigi1103):

help?

Vocaloid (vocaloid):

right... r + 2r = 3r so far we have: 3r + 3 = 75 now subtract 3 from each side, what do you get?

OpenStudy (gigi1103):

these are the answers A) 24 B) 47 C) 48 D) 51

Vocaloid (vocaloid):

3r + 3 = 75 now subtract 3 from each side, what do you get?

OpenStudy (gigi1103):

how do you subtract on two sides?

OpenStudy (gigi1103):

???????????

Vocaloid (vocaloid):

there are two sides of the equation 3r + 3 and 75 we start off with 3r + 3 = 75 now we take away 3 from each side... 3r + 3 - 3 = 75 - 3 can you take it from here?

OpenStudy (gigi1103):

ok so 3-3=0 75-3=2 3r+0=72?

Vocaloid (vocaloid):

yeah! 3r = 72 now divide both sides by 3

OpenStudy (gigi1103):

ok one sec

OpenStudy (gigi1103):

24?

Vocaloid (vocaloid):

right, so r = 24 now, that means we have 24 red cars, now we can find the number of white cars using w = 2r + 3 so w = ?

OpenStudy (gigi1103):

5

OpenStudy (gigi1103):

5w

OpenStudy (gigi1103):

hello?

Vocaloid (vocaloid):

not quite... r = 24, so we replace "r" with "24" so w = 2r + 3 = 2(24) + 3 = ?

OpenStudy (gigi1103):

ok so 27

Vocaloid (vocaloid):

check your math again 2*(24) = 3 = ?

Vocaloid (vocaloid):

* means multiply

OpenStudy (gigi1103):

ok didnt see the times

OpenStudy (gigi1103):

51

OpenStudy (gigi1103):

@Vocaloid

Vocaloid (vocaloid):

right, so w = 51 and 51 is our answer

OpenStudy (gigi1103):

thanks!!

OpenStudy (gigi1103):

can you help with one more?